# Learning Online Belief Prediction for Efficient POMDP Planning in Autonomous Driving

### v1

@@ Abstract
Effective decision-making in autonomous driving relies on accurate inference of other traffic agents’ future behaviors. To achieve this, we propose an online learning-based behavior prediction model and an efficient planner for Partially Observable Markov Decision Processes (POMDPs). We develop a Transformer-based prediction model, enhanced with a recurrent neural memory model, to dynamically update latent belief states and infer the intentions of other agents. The model can also integrate the ego vehicle’s intentions to reflect closed-loop interactions among agents, and it learns from both offline data and online interactions. For planning, we employ an option-based Monte-Carlo Tree Search (MCTS) planner, which reduces computational complexity by searching over action sequences. Inside the MCTS planner, we use predicted long-term multi-modal trajectories to approximate future updates, which eliminates iterative belief updating and improves the running efficiency. Our approach also incorporates deep Q-learning (DQN) as a search prior, which significantly improves the performance of the MCTS planner. Experimental results from simulated environments validate the effectiveness of our proposed method. The online belief update model can significantly enhance the accuracy and temporal consistency of predictions, leading to improved decision-making performance. Employing DQN as a search prior in the MCTS planner considerably boosts its performance and outperforms an imitation learning-based prior. Additionally, we show that the option-based MCTS substantially outperforms the vanilla method in terms of performance and efficiency.
@@
POMDP, Behavior Prediction, Belief Update, Monte-Carlo Tree Search, Interactive planning
@@ I Introduction
Decision-making under uncertainties is crucial for the safety of autonomous driving systems. In particular, the uncertainty in human road users’ behaviors is a primary source of uncertainties [1], which imposes significant challenges to the safe navigation of autonomous vehicles (AVs) in real-world scenarios. The Partially Observable Markov Decision Process (POMDP) [2, 3] offers a mathematically sound framework to address this problem. However, current POMDP planners have limited capabilities, as they only represent the hidden states of other agents with specific semantic meanings and cannot integrate with deep learning-based prediction models. This restricts their capability and scalability in more diverse and complex real-world situations. Therefore, we aim to develop a deep learning-based online behavior prediction model and an efficient planner for POMDPs in autonomous driving.
@@
Figure 1: Illustration of our proposed planning approach. We utilize a neural memory-based belief update model to continually update the human agent’s intentions over time based on new observations and the AV’s actions. An option-based MCTS planner, guided by a learned Q-value function, searches for the optimal action based on current belief states.
@@
One challenge involves adapting an existing open-loop prediction model to an online and closed-loop setting. We propose a neural memory-based belief update model that optimizes closed-loop prediction performance and learns through online interactions. Specifically, our proposed model includes a Transformer-based encoder that maps the observation of an agent (in the context of maps and other agents) to latent space. At each time step in online training or testing, we utilize a gated recurrent unit (GRU) model to update the latent belief state of each tracked agent by considering its last latent belief state, the current latent observation, and the intention of the ego agent. We then employ a decoder that maps the latent belief state back into the explicit belief state, represented by a probabilistic distribution of intentions (long-term trajectories) for downstream planning. Our model can significantly improve the consistency and probability estimation compared to offline prediction models that are designed, trained, and evaluated in an open-loop manner. Our model is also able to capture the dynamic interactions between the AV and other agents, as opposed to open-loop conditional prediction models [4, 5, 6].
@@
Another challenge lies in developing a computationally efficient planner from POMDPs. To address this, we adopt the Monte-Carlo tree search (MCTS) algorithm [7, 8, 9] and incorporate several enhancements to reduce the computational complexity for real-world applicability. First, we leverage an option-based MCTS method that searches over options (i.e., action sequences or motion primitives) [10, 11], allowing a more in-depth search within a limited computation budget. Additionally, we use predicted multi-modal long-term trajectories from the online prediction model to approximate future updates for other agents within the planning horizon, which improves computational efficiency while ensuring planning performance. Lastly, we train a deep Q-value network based on the environment reward feedback during the online learning process, which is used as the search prior for the MCTS planner and significantly improves its planning performance.
The core idea of our approach is illustrated in Fig. 1, and the main contributions of this paper are summarized as follows.
  1. 1.
@@
    We propose a neural memory-based belief update model that aids in online and closed-loop behavior prediction for POMDP planning.
  2. 2.
@@
    We introduce an option-based MCTS planning method, coupled with a Q-value function network that acts as a heuristic guide during the search process.
  3. 3.
@@
    We establish an online learning framework of the belief prediction and Q-value network and validate our proposed method with a large-scale driving dataset and simulated driving environment.
@@ II Related Work
II-A Behavior Prediction
@@
Recent advancements in deep learning-based prediction models have greatly improved accuracy in various benchmarks [12, 13, 14]. However, when applying these models in online and closed-loop settings, temporal consistency becomes a major problem because these models are all designed to work with independent frames and are trained offline. To address this problem, the motion forecasting with self-consistent constraints (MISC) method [15] has proposed self-consistent losses in training. However, MISC still uses independent frames as model inputs and is trained offline. The concept of streaming motion prediction [16] has been proposed recently, where a neural differentiable filter is used to update the predicted states of agents at each timestep. Our method differs from it in that we update the latent belief states of agents, which ensures temporal consistency, improves prediction accuracy, and allows for incorporating the AV’s influence in the update. Moreover, many studies focus on ego-conditioned prediction (i.e., incorporating the future actions of the AV into the predictions of other agents) to make interaction-aware decisions [17, 18, 19]. However, most of these models are trained in an open-loop manner. In contrast, our model is trained online and capable of modeling closed-loop interactions.
II-B POMDP in Autonomous Driving
POMDP is a mathematically principled framework for decision-making under uncertainty, which has been applied in autonomous driving tasks. [20] proposed a POMDP planner for automated driving that considers human internal states, represented by values of the parameters of driving behavior models. [1] used possible routes followed by other vehicles as the hidden states in POMDP, and the belief states can be estimated over time by particle filter. [21] proposed a POMDP planning method by estimating drivers’ intelligence levels and degrees of rationality through probabilistic inference. However, these planners only infer the semantic parameters of some dynamics or probabilistic models, which limits their capability in more complex scenarios with heterogeneous agents. In contrast, our approach employs latent belief states for other agents, which can be recovered to their future trajectories. Another thread of research is to incorporate neural network models in POMDP [22, 23, 24], and one effective approach is recurrent latent (belief) state update [25]. Therefore, we use the recurrent GRU model to track the latent belief states of other agents and iteratively update their belief states while considering the intention of the ego vehicle.
II-C MCTS Planning in Autonomous Driving
MCTS algorithm has achieved great success in solving complex MDP problems such as board games and Atari games [26], and it has been increasingly applied to autonomous driving tasks [7, 8, 27]. The MuZero method [28] has significantly improved the scalability of MCTS by integrating learned models, such as transition dynamics and value functions. Our work focuses on learning the behavior prediction model for other road users to approximate future transitions. Additionally, to overcome the issue of high-dimensional and continuous search space, one practical solution is using an option-based approach that searches over action sequences instead of raw control actions [10], which effectively reduces search depth. Another improvement is to use learned heuristic models to guide the tree search process and find the optimal solution. For instance, [7] and [8] incorporated a trajectory prediction model as heuristics in tree search, improving both efficiency and performance. In our work, we utilize deep Q-learning as a search prior in the MCTS planner, which shows significant improvements in planning performance.
@@ II-C MCTS Planning in Autonomous Driving
Figure 2: Illustration of the proposed POMDP decision-making framework. (a) Diagram of the POMDP planning process. At each time step, we use a Transformer encoder to map the observation into latent space, which is used to update the latent belief state. We use an MLP decoder to project the latent belief state into probabilistic future trajectories. (b) Option-based MCTS planner. The planner takes as input the estimated future states of other agents and searches over options. Additionally, we employ a learned Q-value network to guide the selection step at the root node. (c) Structure of the belief update network. The information of the previous latent belief state, current latent observation, and the ego agent’s action are fused in the update process.
@@ III-A Problem Statement
We formulate the decision-making problem as a POMDP, which is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{Z},\mathcal{T},\mathcal{R})$. $\mathcal{S}$ is the state space and $\mathbf{s}_{t}^{i}$ denotes the state of agent $i$ and time $t$. $\mathbf{a}_{t}^{e}\in\mathcal{A}$ is the action space for the ego agent. $[\mathbf{o}_{t}^{i},\mathbf{o}_{t}^{m}]\in\mathcal{O}$ is the observation space, where $\mathbf{o}^{i}_{t}$ denotes the observed physical state of the agent (e.g., position, orientation, velocity, and size) and $\mathbf{o}^{m}_{t}$ is the shared road map. $\mathcal{Z}(\mathbf{s}_{t}|\mathbf{o}_{t})$ is the observation function, $\mathcal{T}(\mathbf{s}_{t+1}|\mathbf{a}_{t},\mathbf{s}_{t})$ is the state transition function, and $\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})$ is the immediate reward function.
The ego (decision-making) agent cannot directly observe the internal state $\mathbf{s}^{i}$ of other agents, which can encompass attributes such as the agent’s navigational goals, behavioral traits, and intentions [2]. Therefore, the ego agent maintains a belief $\mathbf{b}_{t}=\{\mathbf{b}_{t}^{i}\}_{i=1:N}$ as an estimate of the states of its surrounding $N$ agents. The belief state starts with an initial belief $\mathbf{b}_{0}^{i}$ and is updated $\mathbf{b}_{t}^{i}$ at each time step $t$ according to a belief update function $\rho$:
@@
We assume that the ego agent makes decisions based on its current belief of the surrounding agents; thus, the entire history of observations is taken into consideration implicitly in the ego agent’s actions. We aim to find the optimal ego policy, denoted by $\pi(\mathbf{a}_{t}|\mathbf{b}_{t})$, that maximizes the expected total discounted reward:
@@
Specifically, we represent the state of an agent $\mathbf{s}^{i}$ as its intention or planned future trajectory. We maintain a latent belief state of each agent in our model, which can be decoded back to an explicit belief state, such as probabilistic intentions or multi-modal future trajectories. To achieve this, we employ an encoder network that maps the observation to a latent space, and a GRU network to update the latent belief state for each agent, which is then decoded to possible trajectories. The overall POMDP planning framework is illustrated in Fig. 2.
@@ III-B Neural Recurrent Belief Update
Observation Encoding. The observation space includes the historical trajectories of traffic agents (including the ego) $\mathbf{o}^{a}\in\mathbb{R}^{N_{a}\times T_{h}\times D_{a}}$ and map polylines $\mathbf{o}^{m}\in\mathbb{R}^{N_{m}\times N_{w}\times D_{m}}$, where $N_{a}$ is the number of agents, $T_{f}$ is the fixed history steps, $N_{m}$ is the number of polylines, and $N_{w}$ is the number of waypoints in a polyline. Note that their positional attributes are normalized according to their respective local coordinates. We encode the historical states of agents using GRU networks and multi-layer perceptron (MLP) networks to encode the map elements, and we apply max-pooling to reduce the waypoint axis. Finally, we concatenate these elements to obtain an encoding of the scene $\mathbf{h}\in\mathbb{R}^{(N_{a}+N_{m})\times D}$, where $D$ is the hidden feature dimension.
@@
Belief Update. We denote the belief update network as $f_{\rho}$, which takes as input the last latent belief of the agent $\mathbf{\tilde{b}}_{t-1}^{i}\in\mathbb{R}^{D}$, the current latent observation $\mathbf{z}_{t}^{i}$, as well as the ego agent’s selected option (short-term planned trajectory) relative to the agent $\mathbf{a}_{t-1}^{i,e}\in\mathbb{R}^{T_{o}\times 3}$. We set the initial latent belief as $\mathbf{\tilde{b}}_{0}^{i}=\mathbf{z}_{0}^{i}$, and the belief update process can be formulated as follows.
@@
where $\mathbf{b}^{i}_{t}$ represents the explicit belief state, and $f_{\phi}$ is the trajectory decoder network that decodes multimodal and uncertainty predictions from the latent belief state.
The structure of the belief update network is illustrated in Fig. 2(c). We first account for the influence of the ego agent’s action on the belief of the tracked agent. For this, we use two MLPs - one is utilized to encode the ego agent’s relative trajectory, and the other is used to predict a gating value. We use this gating value to control the flow of information to the belief update because some agents may not be influenced by the ego agent. Afterward, the ego agent’s influence is combined with the latent observation to update the belief state through the GRU cell. Note that some agents may be occluded for a few timesteps, and we take the last available belief state of those agents in the update.
Trajectory Decoding. Given an agent’s latent belief state $\mathbf{\tilde{b}}_{t}^{i}$, we use an MLP $f_{\phi}$ to decode from the latent belief to explicit belief, which is represented by a Gaussian mixture model (GMM) of future trajectories $\mathbf{b}^{i}\in\mathbb{R}^{M\times T_{f}\times D_{f}}$. This GMM comprises $M$ modalities and $T_{f}$ future timesteps, and features at each step include $x$ and $y$ coordinates, variances $\sigma_{x}$ and $\sigma_{y}$, and probability $p$. It is important to note that the decoder $f_{\phi}$ can also be used for the latent observation $\mathbf{z}^{i}$, as in most existing open-loop motion prediction models.
III-C Option-based MCTS Planner
We use the belief states of agents $\mathbf{b}_{t}=\{\mathbf{b}_{t}^{i}\}_{i=1:N}$ (multi-modal long-term trajectories) to approximate the future updates within the planning horizon. This effectively simplifies the POMDP problem into an MDP problem [20], which can be efficiently solved using the MCTS algorithm. We employ receding horizon planning and construct a new policy for the current belief each time a new observation is received. Furthermore, to optimize the quality and efficiency of the MCTS planner, we employ an option-based approach.
Options. An option is a sequence of actions over a time duration [11], and options can be learned from data and can differ in length [10, 29]. For simplicity, we employ fixed-length heuristic options (high-level actions, like overtaking and turning). We assume that the routing module provides a reference path, such that we can generate options based on this reference path. At each node, options are generated using the available actions for acceleration ($a_{s}$) and lateral speed ($v_{l}$), and the resulting states are calculated in the Frenet frame:

### v2

@@ Abstract
Effective decision-making in autonomous driving relies on accurate inference of other traffic agents’ future behaviors. To achieve this, we propose an online belief-update-based behavior prediction model and an efficient planner for Partially Observable Markov Decision Processes (POMDPs). We develop a Transformer-based prediction model, enhanced with a recurrent neural memory model, to dynamically update latent belief state and infer the intentions of other agents. The model can also integrate the ego vehicle’s intentions to reflect closed-loop interactions among agents, and it learns from both offline data and online interactions. For planning, we employ a Monte-Carlo Tree Search (MCTS) planner with macro actions, which reduces computational complexity by searching over temporally extended action steps. Inside the MCTS planner, we use predicted long-term multi-modal trajectories to approximate future updates, which eliminates iterative belief updating and improves the running efficiency. Our approach also incorporates deep Q-learning (DQN) as a search prior, which significantly improves the performance of the MCTS planner. Experimental results from simulated environments validate the effectiveness of our proposed method. The online belief update model can significantly enhance the accuracy and temporal consistency of predictions, leading to improved decision-making performance. Employing DQN as a search prior in the MCTS planner considerably boosts its performance and outperforms an imitation learning-based prior. Additionally, we show that the MCTS planning with macro actions substantially outperforms the vanilla method in terms of performance and efficiency.
@@
POMDP, Behavior Prediction, Belief Update, Multi-agent Modeling, MCTS, Interactive Planning
@@ I Introduction
Decision-making under uncertainties is crucial for the safety of autonomous driving systems. In particular, human traffic participants’ behaviors are a primary source of uncertainties, which imposes significant challenges to the safe navigation of autonomous vehicles (AVs) in real-world scenarios. The Partially Observable Markov Decision Process (POMDP) [1] offers a mathematically sound framework to address this problem. However, most POMDP planners in autonomous driving have limited capabilities, as they only represent the hidden states of other agents with specific semantic meanings [2, 3, 4], which restricts their capability and scalability in complex real-world situations. To overcome this issue, several studies propose enhancing POMDP planning with learned neural network policy and value priors [5, 6, 7]. However, within the context of autonomous driving, learning the state transition function (i.e., predicting the future actions of other agents) is the most challenging part of POMDP planning. Therefore, we aim to integrate POMDP planning with deep learning-based prediction models and develop an online behavior prediction model for effective planning.
@@
Figure 1: Illustration of our proposed planning approach. We utilize a neural memory-based belief update model to continually update other human agents’ intentions over time based on new observations and the AV’s actions. A macro-action-based MCTS planner, guided by a learned Q-value function, searches for the approximately optimal action based on the current belief state.
@@
One challenge is adapting the prediction model to an online and closed-loop setting. We propose a neural memory-based belief update model that optimizes closed-loop prediction performance and learns through online interactions. Specifically, our proposed model uses a Transformer-based encoder to map the observation to latent space. At each time step in online training or testing, we utilize a gated recurrent unit (GRU) model to update the latent belief state of the AV about each tracked agent by considering their last latent states, the current latent observations, and the intention of the ego agent. We then employ a decoder to map the latent belief state back into the explicit belief state, represented by distributions of intentions (long-term trajectories) for other agents. Our neural memory model can significantly improve the temporal consistency and probability estimation accuracy of behavior prediction models in online testing. Our model is designed to capture the dynamic interactions between the AV and other agents, as opposed to open-loop conditional prediction models [8, 9].
@@
Another challenge lies in developing a computationally efficient POMDP planner for AVs. We adopt the Monte-Carlo tree search (MCTS) algorithm [10, 11] and incorporate several enhancements. First, we leverage a macro-action-based method that searches over action sequences or motion primitives [12, 13], allowing a more in-depth search within a limited computation budget. Additionally, we use predicted multi-modal long-term trajectories from the online prediction model to approximate future transitions for other agents within the planning horizon, which improves computational efficiency while ensuring planning performance. Lastly, we train a deep Q-value network based on the environment reward feedback during the online learning process, which is used as the search prior for the MCTS planner.
In summary, we focus on the decision-making under uncertainty problem for AVs, and we have proposed several learning-based enhancements for effective POMDP planning. The core idea of our approach is illustrated in Fig. 1, and the main contributions are outlined below:
  1. C1)
@@
    We propose a neural memory-based belief update model that provides online and closed-loop agent behavior prediction for AV planning. Refer to Section III-B.
  2. C2)
@@
    We introduce a macro-action-based MCTS planning method for AVs, which integrates the online prediction model for future approximation and a Q-value function network that acts as a heuristic guide. See Section III-C.
  3. C3)
@@
    We establish an online learning framework of belief update model and Q-value network and validate our proposed method with a real-world driving dataset and simulated driving environment. Refer to Section III-D.
@@ II Related Work
POMDP and MCTS in Autonomous Driving. POMDP is a mathematically principled framework for decision-making under uncertainty, which is increasingly being used in AV planning. However, conventional methods are limited to handling simple POMDPs with semantic unobservable properties, such as parameters of driving behavior models [2], possible routes to follow [3], cooperation levels [4], or intelligence levels [14]. MCTS algorithms have been identified as an effective approach to solving POMDPs [15], and we convert POMDP to belief-state MDP [5] to solve it more efficiently. Using learned models, including the transition dynamics, policy, and value functions, can significantly improve the performance and scalability of MCTS [16]. [6] and [7] proposed using learned policy and value functions to guide the belief tree search for solving crowd navigation tasks. [4] proposed using neural networks to estimate the values of leaf nodes and guide the tree search for interactive merging scenarios. However, existing models still presume semantic belief parameters in POMDPs and fail to generalize to complicated situations. Our work addresses this gap by modeling the belief state as general behaviors or intentions of other road users and uses learned models to approximate future transitions, which is critical yet challenging in autonomous driving tasks. Although previous studies have incorporated learning-based behavior prediction models into MCTS planning [10, 11], they typically rely on MDP settings without online belief updates, and the prediction models are often trained offline. Our approach develops an online prediction model that utilizes a neural network for online belief updating (C1). Additionally, we leverage a learned deep Q-network as a search heuristic in the MCTS planner, which shows significant improvements in planning performance (C3).
Employing recurrent networks to track and update (latent) belief states [17, 18] is an effective approach for POMDPs. In our AV planning approach, we utilize a recurrent GRU network to continually update the latent belief state concerning the behaviors of other traffic participants, by distributing the ego agent’s belief state across these gents. Furthermore, to deal with a high-dimensional and continuous search space, we implement a macro-action-based method (C2) that searches over action sequences instead of raw control actions [12], which effectively reduces search depth. Although previous studies have proposed learning-based models in POMDPs and neural belief updaters, adapting these methods for driving applications remains challenging. To this end, we propose a systematic framework to enhance POMDP planning for AVs.
@@
Behavior Prediction. Recent advancements in learning-based prediction models have greatly improved accuracy in various benchmarks [19, 20, 21]. However, when applying these models in online and closed-loop settings, prediction consistency becomes a major problem because these models are all designed to work with independent inputs and are trained offline. To address this problem, the motion forecasting with self-consistent constraints (MISC) method [22] introduced self-consistent losses in training, but it still uses independent frames as inputs and is trained offline. The concept of streaming motion prediction [23] has been proposed recently, where a neural differentiable filter is used to update the predicted states of agents at each timestep. Our method updates the latent belief state instead, which ensures performance and temporal consistency and allows for incorporating the AV’s influence in the update. Moreover, many studies focus on ego-conditioned prediction (i.e., incorporating the future actions of the AV into the predictions of other agents) to make interaction-aware decisions [24, 25]. However, most of these models are trained in an open-loop manner. In contrast, our model is trained online and capable of modeling closed-loop interactions (C3).
@@
Figure 2: Illustration of the proposed POMDP decision-making framework. (a) Diagram of the POMDP planning process. At each time step, we use a Transformer encoder to map the observation into latent space, which is used to update the latent belief state. We use an MLP decoder to project the latent belief state into probabilistic future trajectories per agent. (b) Macro-action-based MCTS planner. The planner takes as input the estimated future states of other agents and searches over macro-actions. Additionally, we employ a learned Q-value network to guide the selection step at the root node. (c) Structure of the belief update network. The information of the previous latent belief state, current latent observation, and the ego agent’s action are fused in the update process.
@@ III-A Problem Statement
We formulate the AV decision-making problem as a POMDP, which is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{Z},\mathcal{T},\mathcal{R})$. $\mathcal{S}$ is the state space and $\mathbf{s}_{t}^{i}$ denotes the state of traffic agent $i$ and time $t$. $\mathbf{a}_{t}^{e}\in\mathcal{A}$ is the action space for the ego agent. $[\mathbf{o}_{t}^{i},\mathbf{o}_{t}^{m}]\in\mathcal{O}$ is the observation space, where $\mathbf{o}^{i}_{t}$ denotes the observed physical state of the agent (e.g., position, orientation, velocity, and size) and $\mathbf{o}^{m}_{t}$ is the shared road map. $\mathcal{Z}(\mathbf{o}_{t}|\mathbf{s}_{t})$ is the observation function, $\mathcal{T}(\mathbf{s}_{t+1}|\mathbf{a}_{t},\mathbf{s}_{t})$ is the state transition function, and $\mathcal{R}(\mathbf{s}_{t},\mathbf{a}_{t})$ is the immediate reward function.
The ego (decision-making) agent cannot directly observe the internal state $\mathbf{s}^{i}$ of other agents, which can encompass attributes such as the agent’s navigational goals, behavioral traits, and intentions [1]. Therefore, the ego agent maintains a belief state $\mathbf{b}_{t}=\{\mathbf{b}_{t}^{i}\}_{i=1:N}$ as an estimate of the intentions of its surrounding $N$ agents. The belief state starts with an initial belief $\mathbf{b}_{0}^{i}$ and is updated $\mathbf{b}_{t}^{i}$ at each time step $t$ according to a belief update function $\rho$:
@@
We assume that the ego agent makes decisions based on its current belief of the surrounding agents; thus, the entire history of observations is taken into consideration implicitly in the ego agent’s actions. We aim to find the approximately optimal ego policy, denoted by $\pi(\mathbf{a}_{t}|\mathbf{b}_{t})$, that maximizes the expected total discounted reward:
@@
Specifically, we represent the state of an agent $\mathbf{s}^{i}$ as its intention or future plan. We maintain a latent belief of the ego agent about the states of other agents, which can be decoded back to an explicit belief state, such as probabilistic intentions or multi-modal future trajectories. To achieve this, we employ an encoder network that maps the observation to a latent space, and a GRU network to update the latent belief state, which is then decoded to possible trajectories per agent. The overall POMDP planning framework is illustrated in Fig. 2.
@@ III-B Neural Recurrent Belief Update
Observation. The observation space [19] typically includes the historical trajectories of agents (including the ego) $\mathbf{o}^{a}\in\mathbb{R}^{N_{a}\times T_{h}\times D_{a}}$ and map polylines $\mathbf{o}^{m}\in\mathbb{R}^{N_{m}\times N_{w}\times D_{m}}$, where $N_{a}$ is the number of agents, $T_{f}$ is the fixed history steps, $N_{m}$ is the number of polylines, and $N_{w}$ is the number of waypoints in a polyline. Note that their positional attributes are normalized according to their respective origin coordinates. We encode the historical states of agents using GRU networks and multi-layer perceptron (MLP) networks to encode the map elements, and we apply max-pooling to reduce the waypoint axis. Finally, we concatenate these elements to obtain an encoding of the scene $\mathbf{h}\in\mathbb{R}^{(N_{a}+N_{m})\times D}$, where $D$ is the hidden feature dimension.
@@
Belief Update. We denote the belief update network as $f_{\rho}$, which takes as input the last latent state $\mathbf{\tilde{b}}_{t-1}^{i}\in\mathbb{R}^{D}$, the current encoded observation $\mathbf{z}_{t}^{i}$, and the ego agent’s planned actions (short-term trajectory) relative to each agent $\mathbf{a}_{t-1}^{i,e}\in\mathbb{R}^{T_{o}\times 3}$. We define the initial latent belief as $\mathbf{\tilde{b}}_{0}^{i}\triangleq\mathbf{z}_{0}^{i}$, and the belief update process can be formulated as follows.
@@
where $\mathbf{b}^{i}_{t}$ represents the explicit belief state, and $f_{\phi}$ is the trajectory decoder network that decodes multi-modal predictions from the latent belief state.
The structure of the belief update network is illustrated in Fig. 2(c). We first account for the influence of the ego agent’s action on the state of the tracked agent. For this, we use two MLPs - one is utilized to encode the ego agent’s relative trajectory, and the other is used to predict a gating value. We use this gating value to control the flow of information to the belief update because some agents may not be influenced by the ego agent. Afterward, the ego agent’s influence is combined with the encoded observation to update the belief state through the GRU cell. Note that some agents may be occluded for a few timesteps, and we take the last available belief state of those agents in the update.
Trajectory Prediction. Given the latent belief state about an agent $\mathbf{\tilde{b}}_{t}^{i}$, we use an MLP $f_{\phi}$ to decode from the latent belief to explicit belief, which is represented by a Gaussian mixture model (GMM) of future trajectories $\mathbf{b}^{i}\in\mathbb{R}^{M\times T_{f}\times D_{f}}$. This GMM comprises $M$ modalities and $T_{f}$ future timesteps, and features at each step include $x$ and $y$ coordinates, variances $\sigma_{x}$ and $\sigma_{y}$, and probability $p$. It is important to note that the decoder $f_{\phi}$ can also be used for the encoded observation $\mathbf{z}^{i}$, as in most existing offline motion prediction models.
III-C Macro-action-based MCTS Planner
We use the belief state of the ego agent over other agents $\mathbf{b}_{t}=\{\mathbf{b}_{t}^{i}\}_{i=1:N}$ to approximate the future updates within the planning horizon (multi-modal long-term trajectories). This effectively simplifies the POMDP problem into a belief-state MDP [2], which can be efficiently solved using the MCTS algorithm. We employ receding horizon planning and construct a new policy for the current belief each time a new observation is received. Furthermore, to optimize the quality and efficiency of the MCTS planner, we employ macro-actions.
Macro-actions. A macro-action is a sequence of instantaneous actions over a time duration [26, 13], and they can be learned from data and differ in length [12]. For simplicity, we employ fixed-length heuristic macro-actions (high-level discrete intentions, such as overtaking and turning). We assume that the routing module provides a reference path, such that we can generate macro-actions based on this reference path. At each node, macro-actions are generated using the available actions for acceleration ($a_{s}$) and lateral speed ($v_{l}$), and the resulting states are calculated in the Frenet frame: